OpenStudy (anonymous):

Find the indicated dot product. r = <9, -7, -8>, v = <3, 4, 7>, w = <6, -9, 7> v ⋅ w <-27, 28, 56> <18, -36, 49> 18 31

OpenStudy (xapproachesinfinity):

do you know how to do dot product

OpenStudy (anonymous):

I'm not the best

OpenStudy (anonymous):

3x6 + 4x(-9) + 7x7

OpenStudy (anonymous):

do i just have to simplify this?

myininaya (myininaya):

yah

OpenStudy (xapproachesinfinity):

oh well you know how to do this :) and yes just simplify

OpenStudy (xapproachesinfinity):

dot product gives a scalar (ie a number)

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

can you help me simplify. That was the part i was having trouble with.

OpenStudy (xapproachesinfinity):

okay 3x6+4x(-9)+7x7=18-36+49=-18+49=31

OpenStudy (anonymous):

Thanks :)

OpenStudy (xapproachesinfinity):

welcome!

OpenStudy (zzr0ck3r):

I would go review some basic math if you are not able to figure out 3*6+4*-9+7*7 I only say this because it seems like if you are trying to do things like the dot product, then you should be able to do basic arithmetic.

OpenStudy (xapproachesinfinity):

I would agree with that too :)